OpenStudy (anonymous):

g(x)=3x-1 domain=(-2,-1,0,1,2) find range

OpenStudy (anonymous):

@Conqueror @cwrw238 @thomaster

OpenStudy (anonymous):

@PaulaLovesSchool13

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

I would love to help :)

OpenStudy (anonymous):

thank you I know that Domain mean x and I have to find the Y values

OpenStudy (anonymous):

but when I add it into the equation it doesn't make sense

OpenStudy (anonymous):

Wait, you do?

OpenStudy (anonymous):

Ugh, I hate these thing -_-

OpenStudy (anonymous):

Sorry Sunny! :(

OpenStudy (anonymous):

you take the Domain and you add it into the equation for example g (0)=3*0-1

OpenStudy (anonymous):

Yes I know but when i do them I get them wrong, sorry maybe ask an online math calculator or mathlete :)

OpenStudy (anonymous):

Again, sorry.

OpenStudy (anonymous):

k @math&ing001 @onlinemathcalculator @mathlete

OpenStudy (anonymous):

I'll stink in here in case someone helps since I struggle wit this. O_O Yes, an 8th grade doesn't know how to do this like me :(

OpenStudy (anonymous):

stick* sorry!

OpenStudy (math&ing001):

That's exactly how you do it @sunnycali ! g(0) = 3*0 - 1 g(1) = 3*1 -1 Remember that multiplication comes before addition (PEMDAS)

OpenStudy (anonymous):

Ig(0)=-1 g(1)=2

OpenStudy (math&ing001):

Yes ! Do it for all the values you have on the domaine to find the range.

OpenStudy (anonymous):

is g(2), 5

OpenStudy (math&ing001):

Yap !

OpenStudy (anonymous):

it says it wrong

OpenStudy (math&ing001):

What values did you enter so I can check ?

OpenStudy (anonymous):

here my answer -2,-1 -1,-2 0,-1 1,2 2,5

OpenStudy (math&ing001):

Well 0, 1, and 2 are correct. But -2 and -1 are not.

OpenStudy (anonymous):

can you help with those

OpenStudy (math&ing001):

g(-2) = 3*(-2) - 1 = -6 - 1 = -7 Can you do g(-1) ?

OpenStudy (anonymous):

g(-1)=3*(-1)-1=-2

OpenStudy (anonymous):

-4

OpenStudy (math&ing001):

Yes !